x^2+ 4x + 8= 8x = 29
This is my problem I am having teouble figuring this out! I dont have a clue on how to do it!! if someone can tell me how to do it not just the answer!! THANX A BUNCH!!
Signed,
dnt know how to do this problem girl!
: x^2+ 4x + 8= 8x = 29
This is my problem I am having teouble figuring this out! I dont have a clue on how to do it!! if someone can tell me how to do it not just the answer!! THANX A BUNCH!!
Signed,
dnt know how to do this problem girl!

Answer by sbarbagallo(4)

I'm assuming your first = sign is actually a - sign.
Here you go:
x^(2)+4x+8-8x=29
Since 4x and -8x are like terms, add -8x to 4x to get -4x.
x^(2)-4x+8=29
Since 8 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 8 from both sides.
x^(2)-4x=-8+29
Add 29 to -8 to get 21.
x^(2)-4x=21
To set the left-hand side of the equation equal to 0, move all the expressions to the left-hand side.
x^(2)-4x-21=0
For a polynomial of the form x^(2)+bx+c, find two factors of c (-21) that add up to b (-4). In this problem 3*-7=-21 and 3-7=-4, so insert 3 as the right hand term of one factor and -7 as the right-hand term of the other factor.
(x+3)(x-7)=0
Set each of the factors of the left-hand side of the equation equal to 0.
x+3=0
Set each of the factors of the left-hand side of the equation equal to 0.
x-7=0
Since 3 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 3 from both sides.
x=-3
Since -7 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 7 to both sides.
x=7
The complete solution is the set of the individual solutions.
x=-3,7